OpenStudy (anonymous):

how do you solve 3x-8+2x+2?

OpenStudy (anonymous):

3x-8+2x+2 5x-6 do you not have it equalling anything?

OpenStudy (anonymous):

5x-6

OpenStudy (anonymous):

it says it + 10 , but i dont see how.explain it to me?

OpenStudy (anonymous):

*=10

OpenStudy (anonymous):

so its saying 3x-8+2x=10? right?

OpenStudy (anonymous):

3x-8+2x+2=10 5x-6=10 5x=16 16/5

OpenStudy (anonymous):

thanks